OpenStudy (kittykitty04):

The price of egg rolls at Angela's Restaurant was raised from 2$ to $3. Their Qd fell from 15,000 to 12,000 per week. A. Using the midpt formula, what is the Ed of demand for Eggrolls. Does it make economic sense? B. Assume the restaurant wants to increase it's sale by 10%, by what % should it lower the price?

OpenStudy (mayank):

Ed= (%age change in quantity demanded)/(%age change in price) = \[\frac{ 15000-12000 }{ \frac{ 15000+12000 }{ 2 }} \div \frac{ $3-$2 }{ \frac{ $3+$2 }{ 2} }\] \[= \frac{ 3000 }{ 13500 } \div \frac{ $1 }{ $2.5 }\] \[=0.56\]

OpenStudy (mayank):

It will actually be -0.56 (12000-15000) .